XYZ Co. recently paid a dividend of $2.50. If dividends are expected to grow at a constant rate of 4% and the investor's required rate of return is 10%, what should the stock value be six years from now?

Group of answer choices

$54.83

$52.72

not enough information

$53.89
Answer #1

Stock Value be 6 years from Now =Dividend Year 0*(1+g)^7/(Required Rate-growth) =2.50*(1+4%)^7/(10%-4%) =54.83
(Option a is correct option)